Solve the initial value problem below using the method of Laplace transforms. w - 6 + 9w = 27t +63, w( - 1) = 3, w(-1) = -
Answer #1

Joli Given 163 = 2 = ه - اما _ ما 0 Taring ;ا ا 6 ادا ۹+ ({43L + +اه اها ها ا ا ا د ا کر رہا (9) - وما 13 ( +۹l 3 = 1 + 27as Now [13 م ا م و (3) 2>= لار 3-3 - C [ 34 33 34 at ري)۴} که ce 3 a e خره -12F :1 را 3 + 34 ] ** و (+3 + 3ع ہو یا 44 - يا ]=) segt tz - الا (۷- کے مکرم 2 (2-4 کا لاہ(( کے) 3 tos =AS(176879) +B (6-6879 te(1²38) toon = (A +e) 8+1-6A +1-36 +D)5 +(9A-697 Now in taking inverna L-1, and using 3 we get 3,6 6 3t with = a e witte 423+(1+3+1 +62-667te36 +9+3t-ge qe 3t +24te 7 3+ (2 - 3 се- 2 с — пе za é 3 + 64 - С) е-3 +4 -qe-3_24 е-2 + 3 - 27 е-3 – чес- 3 rin <a 2 е* 49, с.= це?+ 3 6 - | - - 2 3 То, C2Putting there in wit= (2 e²+9) est (1+3+) clé+51) te+9+3+ 3 3t q e +24t e 3t 3t - 311tt) 3617t) 2 e tote tge 36 3 (tt) e te +